Joe has a total of $200 in his two pockets. He takes one-fourth of the money in his left pocket and puts it in his right pocket. He then takes $20 from his left pocket and puts it in his right pocket. If he now has an equal amount of money in each pocket, then how much money did he originally have in his left pocket?

Answer by ankor@dixie-net.com(22740)

Write an expression for each statement.
:
Joe has a total of $200 in his two pockets.
:
Let x = amt in left pocket originally
Then
(200-x) = amt in right pocket originally
:
He takes one-fourth of the money in his left pocket and puts it in his right pocket.
Left:
(x - .25x)
Simplify
.75x
:
Right
(200-x) + .25x
Simplify
200 -.75x
:
He then takes $20 from his left pocket and puts it in his right pocket.
Left:
.75x - 20
:
Right
200 - .75x + 20
Simplify
220 - .75x
:
If he now has an equal amount of money in each pocket,
left = right
.75x - 20 = 220 - .75x
.75x + .75x = 220 + 20
1.5x = 240
x = 240/1.5
x = $160 in his left pocket originally
:
That means he had 200 - 160 = $40 in his right pocket
:
Check for equality
Left 160 - 40 - 20 = 100
Right 40 + 40 + 20 = 100
:
Did this make sense to you?